# Adaptive and robust watermark against model extraction attack

### v1

@@ IV Experiments and Analysis
Figure 4: Cumulative Distribution Function (CDF) of Sentence Watermark Scores: This graph shows the CDF of sentence watermark scores for $M_{I}$, $M_{L}$, and $M_{O}$. It clearly illustrates that the distribution of the Imitation model is significantly distinct from the other two types.
IV-A Construction of Self-Generated Watermark Datasets
As we are the first to explore the self-generated watermarking capability of language models, we build two LLM self-generated watermark datasets. Given the widespread utilization of question-and-answer interactions in language model services, we built the self-generated watermarks dataset based on two publicly available open-ended question-answering datasets, namely Human ChatGPT Comparison Corpus (HC3) [54] and InstructWild (WILD) [55]. HC3 comprises comparison responses from both human experts and ChatGPT across diverse domains, including finance, medicine, law, and psychology. WILD, on the other hand, is an open QA mind strom dataset rich in varied and imaginative instructions. To simulate the process of a model knowledge thief querying victim language models, we sampled 4,000 instances from each of these datasets and employed our method to embed watermarks into the answers provided by the victim models [^2] Our data and models are available at Github after publication. Then, based on the constructed watermark dataset, we set the watermark threshold for the upcoming experiments. As shown in the Table II, we calculate Watermark Scores for human texts in each domain based on the union of watermark words generated on watermarked datasets. This includes analyses of human responses within HC3 and WILD, where they closely match the domain and are most likely to trigger watermark words unintentionally. For each dataset, we selected texts with the top 1% Sentence Watermark Scores to compute their Watermark Score. In our experiments, we usually set threshold $\theta$ at 0.11.
@@ IV-A Construction of Self-Generated Watermark Datasets
TABLE II: Watermark Score on human dataset. We aggregate the sentence watermark scores of the top 1% of sentences in the dataset to statistically analyze the watermark scores
@@ IV-B Model Extraction Attack Simulation
In our study, we employ ChatGPT [62] as the victim model, querying it and applying our method to generate answers containing watermark words. We use the transformer-based [63] model as the imitation model due to its widespread adoption and effectiveness in text generation. We utilized GPT-2 Large-0.7B [64], LLaMA2-7B [65], and Mistral-7B [66] as backbones for our imitation models. These selections represent a diverse set of models: GPT-2 Large as a homologous smaller model and LLaMA2-7B and Mistral-7B as heterogeneous larger models.
@@
TABLE III: Fixed Training Hyperparameters of Imitation Model
IV-C Evaluation
The effectiveness and learnability are gauged using the sentence watermark score or watermark score, as elucidated in our watermark verification method. Robustness is evaluated based on the retention of the watermark despite the aggressiveness of the watermarked data. If the watermark remains detectable post-attack, we consider it robust. Harmlessness is evaluated by the impact of the watermark on the language model’s performance in its intended tasks. We evaluated the harmlessness from two perspectives: firstly, the harmlessness of the watermarked text, where we observed if there’s any decrease in output quality from the victim model ($M_{V}$) after incorporating the watermark mechanism. Secondly, we evaluate harmlessness post-watermark training by comparing the task performance of the imitation model ($M_{I}$) trained with watermarked data against a model ($M_{L}$) trained with non-watermarked data. Specifically, we assess the language model’s QA task performance using BLEU [68] and ROUGE [69] metrics based on human gold standards. The specific calculation formulas of BLEU and ROUGE we used is in appendix VI-C.
@@ IV-D Performance Evaluation
As shown in Fig 3, we evaluated the effectiveness and harmlessness of watermarks on models fine-tuned with watermarked data ($M_{I}$), with non-watermarked data ($M_{L}$), and original base models ($M_{O}$), on two datasets and three foundational bases. The left bar in our representation indicates the effectiveness of the watermark, quantified specifically by the average Sentence Watermark Score of all texts generated by $M_{S}$. The right bars illustrate the non-intrusiveness of watermarks, reflected through task performance metrics.
@@
Firstly, it is evident from the data that the watermark score of the $M_{I}$ is statistically higher than that of the $M_{L}$ and the $M_{O}$. This finding suggests a distinctive difference in $M_{I}$, which has been trained on watermarked data. We repeated 10 times experiments and conducted further detailed contrastive verification KS tests, as Table IV illustrates. We consistently observed that the average p-values of statistical tests between $M_{I}$ and both $M_{O}$ and $M_{L}$ are significantly less than $10^{-12}$. Furthermore, rapid watermark verification methods yielded p-values consistently below $10^{-9}$, confirming our watermark method’s reliability. In Fig 4, we plotted the cumulative probability distribution of watermark scores for the three types of models, further reflecting the significant statistical differences between $M_{I}$ and both $M_{L}$ and $M_{O}$. This result strongly suggests that $M_{I}$ has effectively learned the watermark patterns, showing a marked preference for watermark tokens.
dataset model p-value(fast) p-value(nW) p-value(ori)
HC3 GPT2-Large 1.73E-15 2.31E-13 2.13E-14
HC3 LLaMA2 9.49E-64 2.41E-29 3.27E-14
HC3 Mistral 6.25E-10 2.29E-110 7.12E-179
WILD GPT2-Large 6.23E-68 7.04E-17 4.64E-12
WILD LLaMA2 1.20E-57 9.85E-43 5.72E-13
WILD Mistral 5.64E-18 3.17E-122 5.53E-214
TABLE IV: p-value of the watermark detection on rapid detection and detailed detection.
ROUGE BLEU PPL
HC3 Victim Model w/ 0.149 0.079 9.94
HC3 Victim Model w/o 0.146 0.075 9.86
WILD Victim Model w/ 0.356 0.194 13.37
WILD Victim Model w/o 0.350 0.206 12.78
@@
TABLE V: We measured the task performance of the victim model, ChatGPT, on the HC3 and WILD datasets for question-answering tasks before and after the addition of our watermarking mechanism. ’w/’ represents with our watermark method, while ’w/o’ represents without our watermark method
@@
In terms of the harmlessness of the watermarks, As Table V depicts, for the victim model $M_{V}$, the implementation of our watermarking mechanism shows negligible differences in performance on QA tasks, indicating minimal impact from system-mode self-watermarking prompts. This is attributed to the fact that our prompt-induced method avoids forcefully distorting the original distribution of the language model. As for the imitation model, we observe that the quality of answers produced by the $M_{I}$ is on the same scale as that of the $M_{L}$, and at times, it even surpasses the answer quality of the $M_{L}$, signifying that the training with watermarked data doesn’t impede the model’s capabilities. This subtlety ensures that the model attackers do not easily detect anomalies or even engage in reverse engineering attacks against the watermarks.
@@ IV-E Effectiveness of Different Watermark Detection Strategies
(a) Different watermark strategy for GPT2 backbone
(b) Different watermark strategy for LLaMA2 backbone
(c) Different watermark strategy for Mistral backbone
Figure 5: We tested the average composite average sentence watermark score of three base models—GPT2, LLaMA2, and Mistral—on the HC3 and WILD datasets.
In evaluating watermarks, we adopt three distinct criteria to examine the watermarks produced by the model from various perspectives.
@@
Generally, the size of $WM_{H}$ is the smallest when compared to $WM_{Q}$ or $WM$. Consequently, the ASWS for $WM_{H}$ tends to be relatively lower in watermark detection as shown in the Fig 5.
@@ IV-F Performance Under Varying Training Adequacies
IV-G Comparison with the baseline model
Our comparative method primarily revolves around watermarking methods for language model IP protection. Given our objective to minimally impact the language model’s distribution, we have limited access to the parameters of the target model (ChatGPT). Consequently, our comparison is constrained to the work conducted by He [15] in this particular setting. The baseline method involves embedding watermarks in the output content of the model to be protected through post-processing. Specifically, this method mainly includes synonym replacement in the output content of the model to be protected. Following this, attackers utilize the altered data to train their own models. When compared to models trained on normal data without knowledge of synonym modification rules, models trained with watermarked data tend to exhibit a preference for synonymous with the embedded watermark.
@@ IV-G Comparison with the baseline model
TABLE VI: HC3 example for synonyms as watermarked word will cause some weird generation in some case. The question is ”why the new seasons for a tv show are typically in fall. Would n’t it make sense for some shows to come out in spring or summer so that there ’s less competition for viewership?”
(a) compare with baseline [15] on HC3
(b) compare with baseline [15] on WILD
Figure 7: Compare with baseline: We compared the linguistic quality (indicated by perplexity, or PPL) and task performance (indicated by ROUGE and BLEU scores) of texts provided for users after our method and baseline methods embedded watermarks in the output of the victim model. The baseline method’s rigid synonym substitution resulted in a loss of quality in the watermarked text outputs.
(a) Performance Evaluation of Using GPT2-Large as the Backbone on HC3
(b) Performance Evaluation of Using LLaMA2 as the Backbone on HC3
(c) Performance Evaluation of Using GPT2-Large as the Backbone on WILD
(d) Performance Evaluation of Using LLaMA2 as the Backbone on WILD
(e) Detailed Distribution of SWS using GPT2-Large as the Backbone on HC3
(f) Detailed Distribution of SWS using LLaMA2 as the Backbone on HC3
(g) Detailed Distribution of SWS using GPT2-Large as the Backbone on WILD
(h) Detailed Distribution of SWS using LLaMA2 as the Backbone on WILD
@@
We have compared our approach with the baseline in these aspects: watermark detection, the quality of the protected model’s output, and performance in downstream tasks. The baseline method yields a p-value of less than $10^{-8}$, while our method achieves a p-value of less than $10^{-17}$. Both results are below the $0.05$ threshold, satisfying the criteria for rejecting the null hypothesis. Due to the different watermark detecting algorithms, we evaluate the output quality of the protected model using the baseline method, which necessitates considerable modifications to the output. These assessments were conducted under conditions with comparable p-value magnitudes (around $10^{-10}$). As the Fig 7 shows, our method can provide users with a better linguistic quality response and a higher QA score in downstream tasks. In certain situations, the baseline method can significantly reduce the original output quality of the model, as illustrated in the Table VI. Forcibly replacing synonyms can change the original meaning of sentences, leading to unsuitable expressions. Although CATER [14] attempts to alleviate this semantic inconsistency, this method needs to get access to the model’s conditional token probabilities, which is not allowed in our setting.
@@
Figure 9: Attackers utilize partial watermarked query outputs for Imitation Model Training. The horizontal axis represents the ratio $\eta$ of watermark data used for training the imitation model out of the total watermark data.
@@ IV-H Robustness to Adversarial Attacks
We simulate three adversarial attack that pose challenges to watermark detection in realistic scenarios: user editing attacks, training with only a selection of watermarked data, and training with a mixture of watermarked and unwatermarked data.
@@ IV-H1 Editing Attack
We tested the resilience of our watermarking method against attackers who modify watermarked texts. Specifically, we conducted experiments in four scenarios: random insertion, random deletion, random replacement, and rearranging token order within sentences while maintaining semantic integrity. Details of these attacks are given below:
Random Insertion: This involves copying a random token from the phrase and inserting it at a random location within the same phrase.
Random Deletion: A random token is selected from the phrase and completely removed.
Random Replacement: In this attack, a random token is selected and substituted with another randomly picked token.
Random Sequence change: Two tokens in the sentence are chosen at random and swapped, altering the sequence but not the overall meaning.
@@
IV-H2 Only train a part of watermarked data
@@
Figure 10: Attackers combine non-watermarked data from other sources for training imitation models: In this scenario, attackers use all the watermark data obtained from queries and mix in non-watermarked data from other sources to train imitation models. We controlled the watermark data at a fixed value (400 entries) and mixed in non-watermarked data, varying the watermark data’s proportion $\rho$ to 10%, 25%, 50%, and 75%.
In this mixed training scenario, we prepared six distinct datasets, each containing a varying ratio, $\eta$, of watermarked data relative to the total query responses data, $A$. The values of $\eta$ were set at 0, 0.1, 0.25, 0.5, and 0.75, and we ensured that the size of the total watermark dataset $|A|$, consistently remained at 4000 instances. An $\eta$ value of 0 indicates that the training dataset for the imitation model contained no query response watermarked data, while a $\rho$ value of 1 suggests that all queried watermarked data were used for training the imitation model. The outputs of the suspect model were then evaluated based on our comprehensive knowledge of the query history.
(a) Evaluating BertScore and Sentence Watermark Score Using GPT2-Large as the Backbone on HC3
(b) Evaluating BertScore and Sentence Watermark Score Using GPT2-Large as the Backbone on WILD
(c) Evaluating BertScore and Sentence Watermark Score Using LLaMA2 as the Backbone on HC3
(d) Evaluating BertScore and Sentence Watermark Score Using LLaMA2 as the Backbone on WILD
@@
Even though attackers use only a portion of the watermarked data for training, we conduct watermark detection on all the queries $Q$ known to have been scraped by the victim model’s owner. As depicted in Fig 9, a distinct statistical discrepancy is apparent between models trained with and without watermarked data, even when only 10% of the watermark data are used for imitation model training. This not only demonstrates the robustness of our watermark verification method but also indicates the learnable nature of the watermarked data. It can also be seen that as $\eta$ increases, the average sentence watermark score rises accordingly. This aligns with our intuition that the more watermarked data used for training the imitation model, the easier it is to perform watermark detection.

### v2

@@ IV Experiments and Analysis
Figure 4: Cumulative Distribution Function (CDF) of Sentence Watermark Scores: This graph shows the CDF of 4000 sentence watermark scores of 4000 samples for $M_{I}$, $M_{L}$, and $M_{O}$. It clearly illustrates that the distribution of the Imitation model is significantly distinct from the other two types.
IV-A Self-Generated Watermark Datasets
As we are the first to explore the self-generated watermarking capability of language models, we build two LLM self-generated watermark datasets. Given the widespread utilization of question-and-answer interactions in language model services, we built the self-generated watermarks dataset based on two publicly available open-ended question-answering datasets, namely Human ChatGPT Comparison Corpus (HC3) [10] and InstructWild (WILD) [11]. HC3 comprises comparison responses from both human experts and ChatGPT across diverse domains, including finance, medicine, law, and psychology. WILD, on the other hand, is an open QA mind strom dataset rich in varied and imaginative instructions. To simulate the process of a model knowledge thief querying victim language models, we sampled 4,000 instances from each of these datasets and employed our method to embed watermarks into the answers provided by the victim models [^2] Our data and models are available at https://github.com/amaoku/ModelShield . Then, based on the constructed watermark dataset, we set the watermark threshold for the upcoming experiments. As shown in the Table II, we calculate Watermark Scores for human texts in each domain based on the union of watermark words generated on watermarked datasets. This includes analyses of human responses within HC3 and WILD, where they closely match the domain and are most likely to trigger watermark words unintentionally. For each dataset, we selected texts with the top 1% Sentence Watermark Scores to compute their Watermark Score. In our experiments, we usually set threshold $\theta$ at 0.11.
@@ IV-A Self-Generated Watermark Datasets
TABLE II: Watermark Score on human dataset. We aggregate the sentence watermark scores of the top 1% of sentences in the dataset to statistically analyze the watermark scores.
@@ IV-B Model Extraction Attack Simulation
In our study, we employ ChatGPT [58] as the victim model, querying it and applying our method to generate answers containing watermark words. We use the transformer-based [59] model as the imitation model due to its widespread adoption and effectiveness in text generation. We utilized GPT-2 Large-0.7B [12], Llama2-7B [13], and Mistral-7B [14] as backbones for our imitation models. These selections represent a diverse set of models: GPT-2 Large as a homologous model and LLaMA-7B and Mistral-7B as heterogeneous large language models.
@@
TABLE III: Training hyperparameters of imitation model.
IV-C Metrics
The effectiveness and learnability are gauged using the sentence watermark score or watermark score, as elucidated in our watermark verification method. Robustness is evaluated by whether the watermark can retain its effectiveness after being attacked. Harmlessness is evaluated by the impact of the watermark on the language model’s performance in its intended tasks. We evaluated the harmlessness from two perspectives: firstly, the harmlessness of the watermarked text, where we observed if there was any decrease in output linguistic quality from the victim model ($M_{V}$) after incorporating the watermark mechanism. Secondly, we evaluate harmlessness by comparing the task performance of the imitation model ($M_{I}$) trained with watermarked data against a model ($M_{L}$) trained with non-watermarked data. Specifically, we assess the language model’s QA task performance using BLEU [61] and ROUGE [62] metrics based on human gold standards. The specific calculation formulas of BLEU and ROUGE we used is in supplementary VI-C.
@@ IV-D Performance Evaluation
As shown in Fig 3, we evaluated the effectiveness and harmlessness of watermarks on models fine-tuned with watermarked data ($M_{I}$), with non-watermarked data ($M_{L}$), and original base models ($M_{O}$) using 4000 samples on two datasets and three foundational bases. The left bar in our representation indicates the effectiveness of the watermark, quantified specifically by the average Sentence Watermark Score of all texts generated by $M_{S}$. The right bars illustrate the non-intrusiveness of watermarks, reflected through task performance metrics.
@@
Firstly, it is evident from the data that the watermark score of the $M_{I}$ is statistically higher than that of the $M_{L}$ and the $M_{O}$. This finding suggests a distinctive difference in $M_{I}$, which has been trained on watermarked data. We repeated 10 times experiments and conducted further detailed contrastive verification KS tests, as Table IV illustrates.
We consistently observed that the p-values of statistical tests between $M_{I}$ and both $M_{O}$ and $M_{L}$ are significantly less than $10^{-12}$. In 10 repeated experiments, each with 4000 samples, we did not observe instances where the p-value for the imitation model ($M_{I}$) exceeded 0.05 (false negative) or where the p-values for the backbone model ($M_{O}$) and the legitimate model ($M_{L}$) fell below 0.05 (false positive). Furthermore, rapid watermark verification methods yielded p-values consistently below $10^{-9}$, confirming our watermark method’s reliability.
In Fig 4, we plotted the cumulative probability distribution of watermark scores for the three types of models, further reflecting the significant statistical differences between $M_{I}$ and both $M_{L}$ and $M_{O}$. This result strongly suggests that $M_{I}$ has effectively learned the watermark patterns, showing a marked preference for watermark tokens.
[…]
HC3 [10] Gpt2-Large 1.73e-15 2.13e-14 2.31e-13
Llama2 9.49e-64 3.27e-14 2.41e-29
Mistral 6.25e-10 7.12e-179 2.29e-110
WILD [11] Gpt2-Large 6.23e-68 4.64e-12 7.04e-17
Llama2 1.20e-57 5.72e-13 9.85e-43
Mistral 5.64e-18 5.53e-214 3.17e-122
TABLE IV: Average P-values of Watermark Detection in Rapid and Detailed Contrastive Verification. In Rapid Verification, only the outputs from the suspect model are used. In Detailed Contrastive Verification, the detector has information about the backbone model ($M_{O}$) used for the model extraction attack. This allows for a comparison with the original backbone model $M_{O}$ or legitimate model $M_{L}$.
[…]
HC3 [10] w/ 0.149 0.079 9.94
w/o 0.146 0.075 9.86
WILD [11] w/ 0.356 0.194 13.37
w/o 0.350 0.206 12.78
@@
TABLE V: Task performance of the victim modelon the HC3 and WILD datasets for question-answering tasks before (w/o) and after (w/) the addition of our watermarking mechanism.
@@
As Table V depicts, for the victim model $M_{V}$, the implementation of our watermarking mechanism shows negligible differences in performance on QA tasks, indicating minimal impact from system-mode self-watermarking prompts. This is attributed to the fact that our instruction-guided method avoids forcefully distorting the original distribution of the language model. As for the imitation model, we observe that the quality of answers produced by the $M_{I}$ is on the same scale as that of the $M_{L}$, and at times, it even surpasses the answer quality of the $M_{L}$, signifying that the training with watermarked data doesn’t impede the model’s capabilities. This subtlety ensures that the attackers do not easily detect anomalies or even engage in reverse engineering attacks against the watermarks.
@@ IV-E Effectiveness of Different Watermark Detection Strategies
(a) Different watermark strategy for Gpt2
(b) Different watermark strategy for Llama2
(c) Different watermark strategy for Mistral
Figure 5: Different watermark strategy. We tested the average sentence watermark score of three base models (Gpt2, Llama2, and Mistral) on the HC3 and WILD datasets.
We adopt three distinct criteria to examine the watermarks produced by the model from various perspectives.
@@
Generally, the size of $WM_{H}$ is the smallest when compared to $WM_{Q}$ or $WM$. Consequently, the ASWS for $WM_{H}$ tends to be relatively lower as shown in the Fig 5.
@@ IV-F Performance Under Varying Training Adequacies
IV-G Comparison with baseline
Our comparative method primarily revolves around watermarking methods for language model IP protection. Given our aim to minimally impact the language model’s performance, we have restricted access to token probability manipulation and only receive the output token of the target model (ChatGPT). Consequently, our comparison is constrained to the work conducted by He [7] in this particular setting. The baseline method involves embedding watermarks in the output content of the model to be protected through post-processing. Specifically, this method mainly includes synonym replacement in the output content of the model to be protected. Following this, attackers utilize the altered data to train their own models. When compared to models trained on normal data without knowledge of synonym modification rules, models trained with watermarked data tend to exhibit a preference for synonymous with the embedded watermark.
@@ IV-G Comparison with baseline
TABLE VI: Examples of texts provided for users comparing with baseline on HC3. Synonyms as watermarked word will cause weird generation in some case.
(a) comp. w. baseline [7] on HC3
(b) comp. w. baseline [7] on WILD
Figure 7: Compare with baseline: We compared the linguistic quality (indicated by PPL) and task performance (indicated by ROUGE and BLEU scores) of texts provided for users after our method and baseline methods embedded watermarks in the output of the victim model. The baseline method’s rigid synonym substitution resulted in a loss of quality in the watermarked text outputs.
(a) Performance Evaluation of Using Gpt2-Large on HC3
(b) Performance Evaluation of Using Llama2 on HC3
(c) Performance Evaluation of Using Gpt2-Large on WILD
(d) Performance Evaluation of Using Llama2 on WILD
(e) Detailed Distribution of SWS using Gpt2-Large on HC3
(f) Detailed Distribution of SWS using Llama2 on HC3
(g) Detailed Distribution of SWS using Gpt2-Large on WILD
(h) Detailed Distribution of SWS using Llama2 on WILD
@@
We have compared ModelShield with the baseline in three aspects: watermark detection, the linguistic quality of the protected model’s output, and performance in downstream tasks. As for watermark detection, the baseline method yields a p-value of less than $10^{-8}$ using the default setting, while ModelShield achieves a p-value of less than $10^{-17}$. Both results are below the $0.05$ threshold, satisfying the criteria for rejecting the null hypothesis.
Due to the different watermark detecting algorithms, we evaluate the output’s linguistic quality and performance in downstream tasks of the protected model by contrasting the baseline method at similar p-value magnitudes (around $10^{-8}$). As the Fig 7 shows, our method can provide users with a better linguistic quality response (15.32% and 13.80% PPL $\downarrow$) and a higher QA score in downstream tasks (27.27% and 29.08% ROUGE $\uparrow$). In certain situations, the baseline method can significantly reduce the original output quality of the model, as illustrated in the Table VI. Forcibly replacing synonyms leads to unsuitable expressions in some cases.
@@
Figure 9: Attackers select watermarked data subsets for imitation model training. The horizontal axis represents the ratio $\eta$ of watermark data used for training the imitation model out of the total watermark data.
@@ IV-H Robustness to Adversarial Attacks
We simulate three adversarial attack that pose challenges to watermark detection in realistic scenarios: user editing attacks, selective training on watermarked data subsets, and training with a mixture of watermarked and unwatermarked data.
@@ IV-H1 Editing Attack
We tested the resilience of our watermarking method against attackers who modify watermarked texts. Specifically, we conducted experiments in four scenarios: random insertion, random deletion, random replacement, and rearranging token order within sentences while maintaining semantic consistency (random scramble). Details of these attacks are given below:
1. Random Insertion: This involves inserting a random token at a random location within the same phrase. 2. Random Deletion: A random token is selected from the phrase and completely removed. 3. Random Replacement: In this attack, a random token is selected and substituted with another randomly picked token. 4. Random Scramble: Two tokens in the sentence are chosen at random and swapped, altering the sequence but not the overall meaning.
@@
IV-H2 Selective Training on Watermarked Data Subsets
@@
Figure 10: Attackers combine non-watermarked data from other sources for training imitation models: In this scenario, attackers use all the watermark data obtained from queries and mix in non-watermarked data from other sources to train imitation models. We controlled the watermark data at a fixed value (400 entries) and mixed in non-watermarked data, varying the watermark data’s proportion $\rho$ to 10%, 25%, 50% and 75%.
In this mixed training scenario, we prepared six distinct datasets, each containing a varying ratio $\eta$. The values of $\eta$ were set at 0, 0.1, 0.25, 0.5, and 0.75, and we ensured that the size of the total watermark dataset $|A|$, consistently remained at 4000 instances. The outputs of the suspect model were then evaluated based on our comprehensive knowledge of the query history.
(a) BertScore and Sentence Watermark Score Using Gpt2 on HC3
(b) BertScore and Sentence Watermark Score Using Gpt2 on WILD
(c) BertScore and Sentence Watermark Score Using Llama2 on HC3
(d) BertScore and Sentence Watermark Score Using Llama2 on WILD
@@
Even though attackers use only a portion of the watermarked data for training, we conduct watermark detection on all the queries $Q$ known to have been scraped by the victim model’s owner. As depicted in Fig 9, a distinct statistical discrepancy is apparent between models trained with and without watermarked data (with a p-value of about $10^{-2}$), even when only 10% of the watermark data (400 samples) are used for imitation model training.
This not only demonstrates the robustness of our watermark verification method but also indicates the learnable nature of the watermarked data. It can also be seen that as $\eta$ increases, the average sentence watermark score rises accordingly. This aligns with our intuition that the more watermarked data used for training the imitation model, the easier it is to perform watermark detection.